# MELODY: Robust Semi-Supervised Hybrid Model for Entity-Level Online Anomaly Detection with Multivariate Time Series

### v1

@@ Abstract.
Do, Not, Us, This, Code, Put, the, Correct, Terms, for, Your, Paper
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; May 13–17, 2024; Singaporeprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
@@ 1. Introduction
Existing anomaly detection techniques on MTS are mostly designed for point-level anomaly detection, which is to identify time points that have anomalous observations from their contextual time points in a single entity (Schmidl et al., 2022), as Fig. 1(a) presents. In contrast, entity-level anomaly detection is remarkably different, which is to detect anomalous entities in a stream of entities (e.g. deployments), where each entity emits MTS, as Fig. 1 (b) presents. In particular, entity-level anomaly detection poses four unique challenges.
@@ 2.1. Problem Statement
Remark. It is noteworthy that although the model $f$ needs to check anomalies at every time point for timely detection and rollback, this problem is different from point-level anomaly detection because (1) the model $f$ is shared across entities that have different entity-specific MTSs; and (2) the label y only marks anomalous entities, without marking anomalous time points which are required by the existing (semi-)supervised methods (Jiang et al., 2021; Schmidl et al., 2022; Huang et al., 2022; Chen et al., 2023).
@@ 4.1.1. Featurizer
As illustrated in Fig. 2(b), the OFE has three Featurizers to generate the score $s_{i,j}^{t}$. Each Featurizer is a class. Given a concrete deployment $\mathcal{X}_{i}$, each Featurizer first initializes a set of instances, one per variate, for some (or all) of the variates using their history, e.g., an instance is initialized using the history ${\textbf{x}}_{i,j}$ for variate $j$.
@@ 4.1.3. Feature Aggregator
To address this challenge, and embed different deployments in the same feature space, we propose a feature aggregator. After we obtain the feature $[\hat{s}_{i,j}^{t},\bar{s}_{i,j}^{t}]$ from Eq. (1) for each univariate time series, we aggregates the features over different services for the same metric. Suppose the set of total 22 metrics is =$\mathcal{M}=\{m_{k}\}_{k=1}^{22}$. Taking $\hat{s}_{i,j}^{t}$ as an example, we perform an aggregation for $m_{k}$
@@
Similarly, we can obtain $\bar{z}_{i,j}^{t}$ from $\bar{s}_{i,j}^{t}$, and this step addresses the issue (2). To address issue (1), if a deployment misses a specific metric value $m_{k}$ and Eq. (2) cannot be applied for $m_{k}$, we perform mean-based imputation on $\hat{z}_{i,k}^{t}$ using the training deployments in $\mathcal{X}$. Thus we align all deployments to the same set of 136 features because there are 63 $\hat{z}_{i,k}^{t}$ and 63 $\bar{z}_{i,k}^{t}$ from the 63 featurizer instances in Sec. 4.1.1. Then each deployment $\mathcal{X}_{i}$ is represented by a vector ${\textbf{z}}_{i}\in\mathbb{R}^{d}$ of fixed dimension $d=136$. To address the potentially noisy imputation on missing metrics, in next section, we propose a semi-supervised hybrid model on $\{{\textbf{z}}_{i}\}_{i=1}^{N}$ for selecting valid features and learning meaningful embeddings using the supervision signals.
@@ 4.2.1. Semi-Supervised Deep One-Class Model
Penalizing the distance of embeddings to the center c forces the network to extract those common factors of variation which are most stable within the dataset. Then normal samples tend to be close to the center, while anomalies are further away.
Although DeepSVDD can learn embeddings that are less sensitive to the noises in ${\textbf{z}}_{i}$ from our Feature Aggregator (Sec. 4.1.3), it can not harness the supervised signals for learning embeddings that resolve the ambiguous anomaly definition as discussed in Challenge 3 in Sec. 1. Recent works also indicated even with a small amount of labels, semi-supervised methods could outperform unsupervised methods significantly on anomaly detection (Han et al., 2022).
@@ Appendix A Appendix
A.1. Implementation Details

### v2

@@ Abstract.
Time series, Anomaly detection, Deep learning
copyright: acmcopyrightdoi: XXXXXXX.XXXXXXXconference: ; ; isbn: 978-1-4503-XXXX-X/18/06ccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
@@ 1. Introduction
Existing anomaly detection techniques on MTS are mostly designed for point-level anomaly detection, which is to identify time points that have anomalous observations from their contextual time points in a single entity (Schmidl et al., 2022), such as a server, as Fig. 1(a) presents. Entity-level anomaly detection is remarkably different. It aims to detect anomalous entities in a stream of entities (e.g. deployments), where each entity emits MTS, as Fig. 1 (b) presents. In particular, entity-level anomaly detection poses four unique challenges.
@@ 2.1. Problem Statement
Remark. It is noteworthy that although the model $f$ should check anomalies at new time points for timely detection for at inference time, this problem is different from point-level anomaly detection because (1) the model $f$ is shared across entities with different MTSs; and (2) at training time, the label y only marks anomalous entities, without marking anomalous time points, which are required by the existing (semi-)supervised methods (Jiang et al., 2021; Schmidl et al., 2022; Huang et al., 2022; Chen et al., 2023).
@@ 4.1.1. Featurizer
As illustrated in Fig. 2(b), the OFE has three Featurizers to generate score $s_{i,j}^{t}$. Each Featurizer is a general class. Given a concrete deployment $\mathcal{X}_{i}$, each Featurizer first initializes a set of instances, one per variate using its history, for some (or all) of the variates, e.g., an instance is initialized using the history ${\textbf{x}}_{i,j}$ for variate $j$.
@@ 4.1.3. Feature Aggregator
To address this challenge, and embed different deployments in the same feature space, we propose a feature aggregator. After we obtain the feature $[\hat{s}_{i,j}^{t},\bar{s}_{i,j}^{t}]$ from Eq. (1) for each univariate time series, we aggregates the features over different services for the same metric. Suppose $m_{k}$ is the $k$-th metric. Taking $\hat{s}_{i,j}^{t}$ as an example, we perform an aggregation for $m_{k}$
@@
Similarly, we can obtain $\bar{z}_{i,j}^{t}$ from $\bar{s}_{i,j}^{t}$, and this step addresses the issue (2). To address issue (1), if a deployment misses a specific metric value $m_{k}$ and Eq. (2) cannot be applied for $m_{k}$, we perform mean-based imputation on $\hat{z}_{i,k}^{t}$ using the training deployments in $\mathcal{X}$. Thus we align all deployments to the same set of 134 features (63 $\hat{z}_{i,k}^{t}$, 63 $\bar{z}_{i,k}^{t}$ from the 63 featurizer instances and 8 meta-data features in Sec. 4.1.1). Then each deployment $\mathcal{X}_{i}$ is represented by a vector ${\textbf{z}}_{i}\in\mathbb{R}^{d}$ of fixed dimension $d=134$.
To address noisy imputation and model the correlation of variates in MTS, next, we propose a semi-supervised model on $\{{\textbf{z}}_{i}\}_{i=1}^{N}$ for learning meaningful embeddings using supervision signals.
@@ 4.2.1. Semi-Supervised Deep One-Class Model
Although DeepSVDD can learn embeddings that are less sensitive to the noises in ${\textbf{z}}_{i}$, it can not harness the supervised signals for learning embeddings that resolve the ambiguous anomaly definition (Challenge 3 in Sec. 1). Recent works also indicated even with a small amount of labels, semi-supervised methods could outperform unsupervised methods significantly on anomaly detection (Han et al., 2022).
@@ 4.2.3. Hybrid Model
4.3. Time Complexity Analysis
We analyzed the time complexity of OFE and SemiAD in detail in Appendix A.1. In summary, the time complexity of MELODY for anomaly inference is approximately $O(T)$, which is efficient as $T$ can be fixed as a constant length of historical time series.
@@ Appendix A Appendix
A.1. Time Complexity Analysis
As an online anomaly detection system, the efficiency for online inference is important. For a new deployment, the computational load of MELODY for online anomaly detection includes the computation for featurizer instances (Sec. 4.1.1) and the inference using SemiAD model. In this section, we first analyze the time complexity for initializing each of the featurizers in Sec. 4.1.1 and their score computation, then analyze the time complexity for SemiAD model inference.
Rule-based Featurizer. At the initialization stage, SbF learns the mean $\mu$ and standard deviation $\sigma$ of all $T$ values in the history ${\textbf{x}}_{i,j}$, so its complexity is $O(1)$. TbF sets a predefined threshold $\tau$ with complexity $O(1)$. CbF sets up a sliding window of size $w_{c}$ with complexity is $O(1)$. At inference stage, SbF and TbF check whether $x_{i,j}^{t}$ ($t>T$) is above their thresholds, so the complexity is $O(1)$. Similarly, CbF checks whether $x_{i,j}^{t}$ is a missing value with complexity $O(1)$. Therefore, the complexity of Rule-based Featurizer is $O(1)$.
Algorithm-based Featurizer. At the initialization stage, SubNN segments the length-$T$ history ${\textbf{x}}_{i,j}$ into subsequences with window size $w_{n}$ and stride 1 to form a set $\mathcal{S}$ of length-$w_{n}$ subsequences, so its complexity is $O(T-w_{n})$. MD only sets up a sliding window of size $w_{m}$, so its complexity is $O(1)$. At inference stage, subNN calculates the distance between a length-$w_{n}$ sliding window ending at time step $t$ and its nearest neighbor in $\mathcal{S}$, so the complexity is $O(Tw_{n})$. MD calculates $m_{\text{obs}}=\text{median}([x_{i,j}^{t-w_{m}},...,x_{i,j}^{t-1}])$, and $m_{\text{dif}}=\text{median}([x_{i,j}^{t-w_{m}+1}-x_{i,j}^{t-w_{m}},...,x_{i,j}^{t-1}-x_{i,j}^{t-2}])$ with complexity $O(w_{m})$. Therefore, the complexity of Algorithm-based Featurizer is $O(Tw_{n}+w_{m})$.
Meta-Data Featurizer. The deployment-level meta-data features are static and can be retrieved and appended to rule-based and algorithm-based features with $O(1)$ time complexity.
On top of the featurizers, there is a time pooling layer (Sec. 4.1.2) for addressing the variable duration of different deployments and a feature aggregator (Sec. 4.1.3) for aligning different deployments into the same feature space. For the time pooling layer, both MaxPool and MeanPool in Eq. (1) were updated incrementally in constant time during online inference, and in parallel for different features, so its complexity is $O(1)$. For feature aggregator, the key computational step in Eq. (2) was implemented in parallel for all metrics, and its time complexity for each metric is $O(n_{i})$ for $n_{i}$ variates in the MTS. Additionally, the imputation step in the feature aggregator takes $O(n_{i})$ time for filling missing values at a time step. Therefore, by parallelizing all featurizer instances, the overall time complexity of the OFE module during online inference is $O(Tw_{n}+w_{m}+n_{i})$.
SemiAD Model Inference. The SemiDOC module in SemiAD uses Eq. (7) for inferring anomaly scores, where c and $R$ were pre-computed during training time and stored for model inference. So the key computation is the embedding $\phi({\textbf{z}}_{\text{new}};\boldsymbol{\theta})$, which takes $O(dd_{\text{max}})$ for an MLP encoder with $d_{\text{max}}$ as the maximal dimension of all layers, where $d$ is the dimension of ${\textbf{z}}_{\text{new}}$. In addition, the LightGBM in SemiAD uses $O(d)$ time for inference with a constant number of estimators (Ke et al., 2017). Therefore, the time complexity for both ensemble methods MELODY-M and MELODY-S are $O(dd_{\text{max}})$ at online inference time.
Summary. Integrating the time complexity of the processes for initializing featurizers, online computation of features, and SemiAD model inference, the time complexity of MELODY for online anomaly detection is $O(Tw_{n}+w_{m}+dd_{\text{max}})$, where $T$ is the length of historical time series, $w_{n}$ is the window size used by SubNN, $w_{m}$ is the window size used by MD, $d$ is the dimension of features input to SemiOC, and $d_{\text{max}}$ is the maximal dimension of all layers in the neural network $\phi(\cdot;\boldsymbol{\theta})$. In practice, $w_{n}$, $w_{m}$, and $d_{\text{max}}$ are set as small constants. According to Sec. 4.1.3, $d=134$ is also a small constant. Therefore, the time complexity is approximately $O(T)$. In our experiments, we used 2-day historical time series in minute-wise, so $T=2880$. We set $w_{n}=100$, $w_{m}=100$, and $d_{\text{max}}=128$. According to Sec. 4.1.3, $d=134$. With this setup, the model is efficient and can perform online anomaly detection.
A.2. Implementation Details